OpenStudy (anonymous):

Which line shows the first error in the solution? A. line 2 B. line 3 C. line 4 D. line 5

OpenStudy (anonymous):

3(x-2)=18 (line 1) 3x-6=18 (line 2) 3x-6+6=18+16 (line 3) 3x=24 (line 4) x=21 (line 5)

OpenStudy (anonymous):

@Elsa213 @Jamie182

OpenStudy (anonymous):

@THEHELPER123

OpenStudy (firekat97):

one of the terms is accidentally changed... look carefully

OpenStudy (anonymous):

i think line 2

OpenStudy (anonymous):

plz help

Mehek (mehek14):

it's not line 2

Mehek (mehek14):

look at line 3 on one side, they add 6 on the other side, they add 16

OpenStudy (anonymous):

oh i see so line 3

Mehek (mehek14):

yes line 3 it should be +6 on both sides

OpenStudy (anonymous):

wait im sorry it was a typo it was not 16 it was 6

OpenStudy (anonymous):

3x-6+6=18+6

Mehek (mehek14):

oh then not line 3 then

OpenStudy (anonymous):

@Rushwr she can help she is a good friend of mine

Mehek (mehek14):

it should be line 5 then

Mehek (mehek14):

because 24/3 = 8 not 21

OpenStudy (anonymous):

thx

Mehek (mehek14):

yw ^_^